OpenStudy (anonymous):

A heated piece of metal cools according to the function c(x) = (.5)x − 9, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 2. What will be the temperature of the metal after five hours?

OpenStudy (ikram002p):

well temperature = ffunction of cools + added device of cooling temperature =g(x) +h(x)

OpenStudy (anonymous):

Oh, ok. So the way to solve this question would be [(.5)x-9] + [-x-2]?

OpenStudy (ikram002p):

we dnt solve it , just see the tempreture when x=5

OpenStudy (anonymous):

By doing that I got the answer of \[-\frac{ 1 }{ 2 }x-11\] Though my answer choices are: -7 Celsius 6 Celsius 9 Celsius or 16 Celsius I've been trying to figure out how to solve this question for an hour now but I'm truly not sure what to do.

OpenStudy (anonymous):

Oh ok, by putting x=5 into the equation, I got -13.5.

OpenStudy (ikram002p):

[(.5)x-9] + [-x-2] [(.5)5-9] + [-5-2]=-13.5 correct

OpenStudy (anonymous):

Okay. What should I do next now that I have -13.5?

OpenStudy (ikram002p):

nothing its the answer :D

OpenStudy (anonymous):

AH! I wrote the question in wrong! In reality, the equation would be [(.5)^x-9]+[-x-2]. By putting in x=5, I end up getting 9, which is one of my answers. Thank you very much for your help and sorry for the confusion!

OpenStudy (ikram002p):

no its ok :D good luck